# Drift2Matrix: Kernel-Induced Self Representation for Concept Drift Adaptation in Co-evolving Time Series

### v2

@@ C.3 Optimization of Nonconvex Problem
The optimization problem of Eq. 3 can be solved by the Augmented Lagrange method with Alternating Direction Minimization strategy Lin et al. (2011). Normally, we require $\mathbf{Z}$ in Eq. 3 to be nonnegative and symmetric, which are necessary for defining the block diagonal regularizer. However, the restrictions on $\mathbf{Z}$ will limit its representation capability. Thus, we introducing an intermediate-term $\mathbf{V}$ and transform Eq. 3 to:
@@
$\begin{split}&\min_{\mathbf{Z,V}}\frac{1}{2}\mathrm{Tr}(\bm{\mathcal{K}}+\mathbf{V}^{\mathrm{T}}\bm{\mathcal{K}}\mathbf{V})-\alpha\mathrm{Tr}(\bm{\mathcal{K}}\mathbf{V})+\frac{\beta}{2}||\mathbf{V-Z}||^{2}+\gamma||\mathbf{Z}||_{\boxed{\scriptstyle k}},\\ &=\min_{\mathbf{Z,V}}\frac{1}{2}||\mathrm{\Phi}(\mathbf{S})-\frac{\alpha}{2}\mathrm{\Phi}(\mathbf{S})\mathbf{V}||^{2}+\frac{\beta}{2}||\mathbf{V-Z}||^{2}+\gamma||\mathbf{Z}||_{\boxed{\scriptstyle k}}\\ &\quad\quad\quad s.t.\ \mathbf{Z}=\mathbf{Z}^{\mathrm{T}}\geq 0,\mathrm{diag}(\mathbf{Z})=0,\mathbf{{1}^{\mathrm{T}}\mathbf{Z}=\mathbf{1}^{\mathrm{T}}}\end{split}$ (15)
@@
Consider that $||\mathbf{Z}||_{\boxed{\scriptstyle k}}=\sum_{i=N-k+1}^{N}\lambda_{i}(\mathbf{L_{\mathbf{Z}}})$ is a nonconvex term. Drawing from the eigenvalue summation property presented in Dattorro (2010), we reformulate it as $\sum_{i=N-k+1}^{N}\lambda_{i}(\mathbf{L_{\mathbf{Z}}})=\min_{\mathbf{W}}<\mathbf{L_{\mathbf{Z}}},\mathbf{W}>$, where $0\preceq\mathbf{W}\preceq\mathbf{I},\mathrm{Tr}(\mathbf{W})=\textit{k}$, see Appendix A for detail. So Eq. 15 is equivalent to
$\min_{\mathbf{Z},\mathbf{V},\mathbf{W}}\frac{1}{2}||\mathrm{\Phi}(\mathbf{S})-\frac{\alpha}{2}\mathrm{\Phi}(\mathbf{S})\mathbf{V}||^{2}+\frac{\beta}{2}||\mathbf{V}-\mathbf{Z}||^{2}$ (16)
$\quad\quad\quad\quad+\gamma<\mathrm{Diag}(\mathbf{Z1})-\mathbf{Z},\mathbf{W}>$
@@
$\{\mathbf{W}^{i+1},\mathbf{V}^{i+1}\}=$ $\arg\min_{\mathbf{W},\mathbf{V}}\frac{1}{2}||\mathrm{\Phi}(\mathbf{S})-\frac{\alpha}{2}\mathrm{\Phi}(\mathbf{S})\mathbf{V}||^{2}$
$\quad\quad\quad+\frac{\beta}{2}||\mathbf{V}-\mathbf{Z}||^{2}$
$\quad\quad\quad+\gamma<\mathrm{Diag}(\mathbf{Z1})-\mathbf{Z},\mathbf{W}>$
@@
$\mathbf{W}^{i+1}=$ $\arg\min_{\mathbf{W}}<\mathrm{Diag}(\mathbf{Z1})-\mathbf{Z},\mathbf{W}>,$ (17)
@@
$\mathbf{Z}^{i+1}=\arg\min_{\mathbf{Z}}\frac{\beta}{2}$ $||\mathbf{V}-\mathbf{Z}||^{2}+\gamma<\mathrm{Diag}(\mathbf{Z1})-\mathbf{Z},\mathbf{W}>$ (19)
$s.t.\ \mathbf{Z}=\mathbf{Z}^{\mathrm{T}}\geq 0,\mathrm{diag}(\mathbf{Z})=0$
@@
$\mathbf{V}^{i+1}$ $=(\mathrm{\Phi}(\mathbf{S})^{\top}\mathrm{\Phi}(\mathbf{S})+\beta\mathbf{I})^{-1}(\alpha\mathrm{\Phi}(\mathbf{S})^{\top}\mathrm{\Phi}(\mathbf{S})+\beta\mathbf{Z})$ (20)
$=(\bm{\mathcal{K}}+\beta\mathbf{I})^{-1}(\alpha\bm{\mathcal{K}}+\beta\mathbf{Z})$
@@
$\mathbf{Z}^{i+1}=\arg\min_{\mathbf{Z}}$ $\frac{1}{2}||\mathbf{Z}-\mathbf{V}+\frac{\gamma}{\beta}(\mathrm{diag}(\mathbf{W})\mathbf{1}^{\mathrm{T}}-\mathbf{W})||^{2}$ (21)
$s.t.\mathbf{Z}=\mathbf{Z}^{\mathrm{T}}\geq 0,\mathrm{diag}(\mathbf{Z})=0$
@@
Proposition C.5
Consider the matrix $\mathbf{A}\in\mathbb{R}^{n\times n}$. Let’s denote $\hat{\mathbf{A}}=\mathbf{A}-$ $\operatorname{Diag}(\operatorname{diag}(\mathbf{A}))$. With this definition, the solution to the following optimization problem:
$\min_{\mathbf{Z}}\frac{1}{2}\|\mathbf{Z}-\mathbf{A}\|^{2},s.t.\ \operatorname{diag}(\mathbf{Z})=0,\mathbf{Z}\geq 0,\mathbf{Z}=\mathbf{Z}^{\top}\text{, }$ (22)
@@
Proof C.6
It is evident that problem Eq. 22 is equivalent to
$\min_{\mathbf{Z}}\frac{1}{2}\|\mathbf{Z}-\hat{\mathbf{A}}\|^{2},s.t.\ \mathbf{Z}\geq 0,\mathbf{Z}=\mathbf{Z}^{\top}.$ (23)
The constraint $\mathbf{Z}=\mathbf{Z}^{\top}$ suggests that $\|\mathbf{Z}-\hat{\mathbf{A}}\|^{2}=\left\|\mathbf{Z}-\hat{\mathbf{A}}^{\top}\right\|^{2}$.
Thus
$\frac{1}{2}\|\mathbf{Z}-\hat{\mathbf{A}}\|^{2}$ $=\frac{1}{4}\|\mathbf{Z}-\hat{\mathbf{A}}\|^{2}+\frac{1}{4}\left\|\mathbf{Z}-\hat{\mathbf{A}}^{\top}\right\|^{2}$
$=\frac{1}{2}\left\|\mathbf{Z}-\frac{\hat{\mathbf{A}}+\hat{\mathbf{A}}^{\top}}{2}\right\|^{2}+c(\hat{\mathbf{A}})$
where $c(\hat{\mathbf{A}})$ only depends on $\hat{\mathbf{A}}$. Hence Eq. 23 is equivalent to
$\min_{\mathbf{Z}}\frac{1}{2}\left\|\mathbf{Z}-\left(\hat{\mathbf{A}}+\hat{\mathbf{A}}^{\top}\right)/2\right\|^{2},s.t.\ \mathbf{Z}\geq 0,\mathbf{Z}=\mathbf{Z}^{\top},$
which has the solution $\mathbf{Z}^{*}=\left[\left(\hat{\mathbf{A}}+\hat{\mathbf{A}}^{\top}\right)/2\right]_{+}$.
@@ Appendix D Domain Agnostic Window Size Selection
Inspired by Bouguessa & Wang (2008), we suggest employing the Minimum Description Length (MDL) principle to determine the optimal window size within the available range. The core concept of the MDL principle revolves around encoding input data based on a specific model, with the aim of choosing the encoding that yields the shortest code length Rissanen (1998). Let WS be the set of all $WS(w)$ values for each window size in the available range. The MDL-selection strategy employed in our work bears resemblance to the MDL-pruning method described in Agrawal et al. (2005); Bouguessa & Wang (2008), where the data is split into two subsets (sparse and dense subsets), with one subset being discarded. In our case, we aim to divide WS into two groups $E$ and $F$. Here, $E$ encompasses the greater values of WS, while the group $F$ encompasses the lower values. Following that, the border separating the two groups is chosen in order to get the optimal window size that minimizes the Minimum Description Length (MDL) criteria. The objective function can be defined according to the MDL criteria:
$J(w)=\min_{w}$ $\log_{2}(\mu_{E})+\sum_{WS(w)\in E}\log_{2}(|WS(w)-\mu_{E}|)$ (25)
$+\log_{2}(\mu_{F})+\sum_{WS(w)\in F}(|WS(w)-\mu_{F}|)$
@@
Appendix E Algorithm of Drift2Matrix
Algorithm 1 Drift2Matrix: Nonlinear concept identification
@@ Appendix E Algorithm of Drift2Matrix
In this section, we delve into the detailed implementation of the Drift2Matrix algorithm, an approach designed for nonlinear concept identification and forecasting in co-evolving time series. The Drift2Matrix algorithm operates in two distinct phases, each encapsulated in its own algorithmic structure.
Algorithm 2 Drift2Matrix: Forecasting Concept and Series Values

### v3

@@ C.1 Optimization of Nonconvex Problem
The optimization problem of Eq. 3 can be solved by the Augmented Lagrange method with Alternating Direction Minimization strategy (Lin et al., 2011). Normally, we require $\mathbf{Z}$ in Eq. 3 to be nonnegative and symmetric, which are necessary for defining the block diagonal regularizer. However, the restrictions on $\mathbf{Z}$ will limit its representation capability. Thus, we introducing an intermediate-term $\mathbf{V}$ and transform Eq. 3 to:
@@
$\begin{split}&\min_{\mathbf{Z,V}}\frac{1}{2}\mathrm{Tr}(\boldsymbol{\mathcal{K}}+\mathbf{V}^{\mathrm{T}}\boldsymbol{\mathcal{K}}\mathbf{V})-\alpha\mathrm{Tr}(\boldsymbol{\mathcal{K}}\mathbf{V})+\frac{\beta}{2}||\mathbf{V-Z}||^{2}+\gamma||\mathbf{Z}||_{\boxed{\scriptstyle k}},\\ &=\min_{\mathbf{Z,V}}\frac{1}{2}||\mathrm{\Phi}(\mathbf{S})-\frac{\alpha}{2}\mathrm{\Phi}(\mathbf{S})\mathbf{V}||^{2}+\frac{\beta}{2}||\mathbf{V-Z}||^{2}+\gamma||\mathbf{Z}||_{\boxed{\scriptstyle k}}\\ &\quad\quad\quad s.t.\ \mathbf{Z}=\mathbf{Z}^{\mathrm{T}}\geq 0,\mathrm{diag}(\mathbf{Z})=0,\mathbf{{1}^{\mathrm{T}}\mathbf{Z}=\mathbf{1}^{\mathrm{T}}}\end{split}$ (10)
@@
Consider that $||\mathbf{Z}||_{\boxed{\scriptstyle k}}=\sum_{i=N-k+1}^{N}\lambda_{i}(\mathbf{L_{\mathbf{Z}}})$ is a nonconvex term. Drawing from the eigenvalue summation property presented in (Dattorro, 2010; Xu et al., 2025), we reformulate it as $\sum_{i=N-k+1}^{N}\lambda_{i}(\mathbf{L_{\mathbf{Z}}})=\min_{\mathbf{W}}<\mathbf{L_{\mathbf{Z}}},\mathbf{W}>$, where $0\preceq\mathbf{W}\preceq\mathbf{I},\mathrm{Tr}(\mathbf{W})=\textit{k}$. So Eq. 10 is equivalent to
$\min_{\mathbf{Z},\mathbf{V},\mathbf{W}}$ $\frac{1}{2}||\mathrm{\Phi}(\mathbf{S})-\frac{\alpha}{2}\mathrm{\Phi}(\mathbf{S})\mathbf{V}||^{2}+\frac{\beta}{2}||\mathbf{V}-\mathbf{Z}||^{2}+\gamma<\mathrm{Diag}(\mathbf{Z1})-\mathbf{Z},\mathbf{W}>$ (11)
@@
$\{\mathbf{W}^{i+1},\mathbf{V}^{i+1}\}=\arg\min_{\mathbf{W},\mathbf{V}}\frac{1}{2}||\mathrm{\Phi}(\mathbf{S})$ $-\frac{\alpha}{2}\mathrm{\Phi}(\mathbf{S})\mathbf{V}||^{2}+\frac{\beta}{2}||\mathbf{V}-\mathbf{Z}||^{2}+\gamma<\mathrm{Diag}(\mathbf{Z1})-\mathbf{Z},\mathbf{W}>$
@@
$\mathbf{W}^{i+1}=\arg\min_{\mathbf{W}}<\mathrm{Diag}(\mathbf{Z1})-\mathbf{Z},\mathbf{W}>,\ s.t.\ 0\preceq\mathbf{W}\preceq\mathbf{I},\mathrm{Tr}(\mathbf{W})=\textit{k}$ (12)
@@
$\mathbf{Z}^{i+1}=\arg\min_{\mathbf{Z}}\frac{\beta}{2}||\mathbf{V}-\mathbf{Z}||^{2}+\gamma<\mathrm{Diag}(\mathbf{Z1})-\mathbf{Z},\mathbf{W}>\ s.t.\ \mathbf{Z}=\mathbf{Z}^{\mathrm{T}}\geq 0,\mathrm{diag}(\mathbf{Z})=0$ (14)
@@
$\mathbf{V}^{i+1}=(\mathrm{\Phi}(\mathbf{S})^{\top}\mathrm{\Phi}(\mathbf{S})+\beta\mathbf{I})^{-1}(\alpha\mathrm{\Phi}(\mathbf{S})^{\top}\mathrm{\Phi}(\mathbf{S})+\beta\mathbf{Z})=(\boldsymbol{\mathcal{K}}+\beta\mathbf{I})^{-1}(\alpha\boldsymbol{\mathcal{K}}+\beta\mathbf{Z})$ (15)
@@
$\mathbf{Z}^{i+1}=\arg\min_{\mathbf{Z}}\frac{1}{2}||\mathbf{Z}-\mathbf{V}+\frac{\gamma}{\beta}(\mathrm{diag}(\mathbf{W})\mathbf{1}^{\mathrm{T}}-\mathbf{W})||^{2}\ s.t.\mathbf{Z}=\mathbf{Z}^{\mathrm{T}}\geq 0,\mathrm{diag}(\mathbf{Z})=0$ (16)
@@
Proposition C.1.
Consider the matrix $\mathbf{A}\in\mathbb{R}^{n\times n}$. Let’s denote $\hat{\mathbf{A}}=\mathbf{A}-$ $\operatorname{Diag}(\operatorname{diag}(\mathbf{A}))$. With this definition, the solution to the following optimization problem: $\min_{\mathbf{Z}}\frac{1}{2}\|\mathbf{Z}-\mathbf{A}\|^{2},s.t.\ \operatorname{diag}(\mathbf{Z})=0,\mathbf{Z}\geq 0,\mathbf{Z}=\mathbf{Z}^{\top}\text{, }$ (17)
@@
Proof.
It is evident that problem Eq. 17 is equivalent to $\min_{\mathbf{Z}}\frac{1}{2}\|\mathbf{Z}-\hat{\mathbf{A}}\|^{2},s.t.\ \mathbf{Z}\geq 0,\mathbf{Z}=\mathbf{Z}^{\top}.$ (18) The constraint $\mathbf{Z}=\mathbf{Z}^{\top}$ suggests that $\|\mathbf{Z}-\hat{\mathbf{A}}\|^{2}=\left\|\mathbf{Z}-\hat{\mathbf{A}}^{\top}\right\|^{2}$. Thus $\frac{1}{2}\|\mathbf{Z}-\hat{\mathbf{A}}\|^{2}$ $=\frac{1}{4}\|\mathbf{Z}-\hat{\mathbf{A}}\|^{2}+\frac{1}{4}\left\|\mathbf{Z}-\hat{\mathbf{A}}^{\top}\right\|^{2}$ $=\frac{1}{2}\left\|\mathbf{Z}-\frac{\hat{\mathbf{A}}+\hat{\mathbf{A}}^{\top}}{2}\right\|^{2}+c(\hat{\mathbf{A}})$ where $c(\hat{\mathbf{A}})$ only depends on $\hat{\mathbf{A}}$. Hence Eq. 18 is equivalent to $\min_{\mathbf{Z}}\frac{1}{2}\left\|\mathbf{Z}-\left(\hat{\mathbf{A}}+\hat{\mathbf{A}}^{\top}\right)/2\right\|^{2},s.t.\ \mathbf{Z}\geq 0,\mathbf{Z}=\mathbf{Z}^{\top},$ which has the solution $\mathbf{Z}^{*}=\left[\left(\hat{\mathbf{A}}+\hat{\mathbf{A}}^{\top}\right)/2\right]_{+}$. ∎
C.2 Permutation Invariance of Representation Matrix in CORAL
Theorem C.2.
In CORAL, the representation matrix obtained for a permuted input data is equivalent to the permutation-transformed original representation matrix. Specifically, let $\mathbf{Z}$ be feasible to $\Phi(\mathbf{S})=\Phi(\mathbf{S})\mathbf{Z}$, then $\tilde{\mathbf{Z}}=\mathbf{P^{T}ZP}$ is feasible to $\Phi(\tilde{\mathbf{S}})=\Phi(\tilde{\mathbf{S}})\tilde{\mathbf{Z}}$.
Proof.
Given a permutation matrix $\mathbf{P}$, consider the self-representation matrix $\tilde{\mathbf{Z}}$ for the permuted data matrix $\mathbf{S}\mathbf{P}$. The objective for $\mathbf{S}\mathbf{P}$ becomes: $\begin{split}\min_{\tilde{\mathbf{Z}}}\left\|\Phi(\mathbf{SP})-\Phi(\mathbf{SP})\tilde{\mathbf{Z}}\right\|^{2}+\Omega(\tilde{\mathbf{Z}}),\quad\text{s.t.}\ \tilde{\mathbf{Z}}=\tilde{\mathbf{Z}}^{\mathrm{T}}\geq 0,\ \text{diag}(\tilde{\mathbf{Z}})=0\end{split}$ (19) By the properties of kernel functions and permutation matrices, we have $\Phi(\mathbf{SP})=\Phi(\mathbf{S}P)$. Substituting this into the objective function for $\tilde{\mathbf{Z}}$, we have: $\begin{split}\min_{\tilde{\mathbf{Z}}}\left\|\Phi(\mathbf{S}\mathbf{P})-\Phi(\mathbf{S})\mathbf{P}\tilde{\mathbf{Z}}\right\|^{2}+\Omega(\tilde{\mathbf{Z}})\quad\text{s.t.}\ \tilde{\mathbf{Z}}=\tilde{\mathbf{Z}}^{\mathrm{T}}\geq 0,\ \text{diag}(\tilde{\mathbf{Z}})=0\end{split}$ (20) Since $P$ is a permutation matrix,$\mathbf{P}\mathbf{P}^{\mathrm{T}}=\mathbf{I}$, the identity matrix. We apply the transformation $\mathbf{P\tilde{Z}P^{\mathrm{T}}}$ to the objective function: $\begin{split}\min_{\tilde{\mathbf{Z}}}\left\|\Phi(\mathbf{S})-\Phi(\mathbf{S})\mathbf{P\tilde{Z}P^{\mathrm{T}}}\right\|^{2}+\Omega(\tilde{\mathbf{Z}})\quad\text{s.t.}\ \tilde{\mathbf{Z}}=\tilde{\mathbf{Z}}^{\mathrm{T}}\geq 0,\ \text{diag}(\tilde{\mathbf{Z}})=0\end{split}$ (21) For the function to be minimized, $\mathbf{P\tilde{Z}P^{\mathrm{T}}}$ must be the optimal representation matrix for $\mathbf{S}$, which is $\mathbf{Z}$. Therefore, $\mathbf{P\tilde{Z}P^{\mathrm{T}}}=\mathbf{Z}$, or equivalently, $\tilde{\mathbf{Z}}=\mathbf{P}^{\mathrm{T}}\mathbf{ZP}$. This result shows that the self-representation matrix $\mathbf{Z}$ for $\mathbf{S}$ transforms to $\mathbf{P}^{\mathrm{T}}\mathbf{ZP}$ for the permuted data matrix $\mathbf{SP}$, demonstrating the invariance of the representation matrix under permutations of the data matrix.
C.3 Manifold Structure Preservation in CORAL
Theorem C.3.
CORAL reveals nonlinear relationships among time series in a high-dimensional space while simultaneously preserving the local manifold structure of series.
Proof.
Optimization problem Eq. 3 can be converted to the form of a matrix trace: $\min_{\mathbf{Z}}\frac{1}{2}\mathrm{Tr}(\boldsymbol{\mathcal{K}}-2\boldsymbol{\mathcal{K}}\mathbf{Z}+\mathbf{Z}^{\mathrm{T}}\boldsymbol{\mathcal{K}}\mathbf{Z})+\gamma||\mathbf{Z}||_{\boxed{\scriptstyle k}},$ (22) In the above, the negative term $-\mathrm{Tr}(\boldsymbol{\mathcal{K}}\mathbf{Z})$ can be transformed into $\min_{\mathbf{Z}}-\mathrm{Tr}(\boldsymbol{\mathcal{K}}\mathbf{Z})$ $=\min_{\mathbf{Z}}\sum_{i=1}^{N}\sum_{j=1}^{N}-\mathrm{\Phi}(\mathbf{S}_{i})^{\mathrm{T}}\mathrm{\Phi}(\mathbf{S}_{j})\mathbf{Z}_{ij}=\min_{\mathbf{Z}}\sum_{i=1}^{N}\sum_{j=1}^{N}-\boldsymbol{\mathcal{K}}(\mathbf{S}_{i},\mathbf{S}_{j})\mathbf{Z}_{ij}$ (23) where $\boldsymbol{\mathcal{K}}(\mathbf{S}_{i},\mathbf{S}_{j})$ indicates the similarity between $\mathbf{S}_{i}$ and $\mathbf{S}_{j}$ in kernel space. It can be seen from Eq. 23 that a large similarity (small distance) $\boldsymbol{\mathcal{K}}(\mathbf{S}_{i},\mathbf{S}_{j})$ tends to cause a large $\mathbf{Z}_{ij}$, and vice versa. This is in fact an kernel extension of preserving local manifold structure in linear space, i.e., $\min_{\mathbf{Z}}\sum_{i=1}^{N}\sum_{j=1}^{N}||\mathbf{X}_{i}-\mathbf{X}_{j}||^{2}\mathbf{Z}_{ij}$ (Nie et al., 2014; Zhan et al., 2018). Suppose a small weight is given to this negative term, which means that the self-representation of the data will take into account the contribution of all other data. Conversely, it will only consider the contribution of other data that are nearest neighbours to the data, thus further enhancing the sparsity of the self-representation $\mathbf{Z}$ while maintaining the local manifold structure. ∎
@@ Appendix D Domain Agnostic Window Size Selection
Inspired by (Bouguessa & Wang, 2008), we suggest employing the Minimum Description Length (MDL) principle to determine the optimal window size within the available range. The core concept of the MDL principle revolves around encoding input data based on a specific model, with the aim of choosing the encoding that yields the shortest code length (Rissanen, 1998). Let WS be the set of all $WS(w)$ values for each window size in the available range. The MDL-selection strategy employed in our work bears resemblance to the MDL-pruning method described in (Agrawal et al., 2005; Bouguessa & Wang, 2008), where the data is split into two subsets (sparse and dense subsets), with one subset being discarded. In our case, we aim to divide WS into two groups $E$ and $F$. Here, $E$ encompasses the greater values of WS, while the group $F$ encompasses the lower values. Following that, the border separating the two groups is chosen in order to get the optimal window size that minimizes the Minimum Description Length (MDL) criteria. The objective function can be defined according to the MDL criteria:
$J(w)=\min_{w}\log_{2}(\mu_{E})+\sum_{WS(w)\in E}\log_{2}(|WS(w)-\mu_{E}|)+\log_{2}(\mu_{F})+\sum_{WS(w)\in F}(|WS(w)-\mu_{F}|)$ (25)
@@
Appendix E Algorithm of CORAL
Algorithm 1 CORAL: Nonlinear concept identification
@@ Appendix E Algorithm of CORAL
Algorithm 2 CORAL: Forecasting Concept and Series Values
@@
In this section, we delve into the detailed implementation of the CORAL algorithm, an approach designed for nonlinear concept identification and forecasting in co-evolving time series. The CORAL algorithm operates in two distinct phases, each encapsulated in its own algorithmic structure.